OpenStudy (anonymous):

Identify the vertex for the graph of y = −3x2 − 6x + 4. (1, 13) (−1, 13) (1, 7) (−1, 7)

OpenStudy (anonymous):

Is it C?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

Then it has to be A because I know X is 1

OpenStudy (anonymous):

first coordinate of the vertex of \(y=ax^2+bx+c\) is \(-\frac{b}{2a}\) and so \(y = -3 x^2-6 x+4\) is \(-\frac{-6}{2\times (-3)}=1\)

OpenStudy (anonymous):

oops i meant \(-1\)! !!

OpenStudy (anonymous):

\[-\frac{-6}{2\times (-3)}=-1\]

OpenStudy (anonymous):

so what is Y?